OpenStudy (anonymous):

Find the derivative of f(x)=x^3+4x-2 using the derivative formula

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

f'=3x^2 +4

OpenStudy (anonymous):

Well, use the difference quotient which is: \[\lim_{h \rightarrow 0}\frac{f(x)-f(x+h)}{h}\]. So: \[\lim_{h \rightarrow 0}\frac{(x^3+4x-2)-((x+h)^3+4(x+h)-2)}{h}\]. Then simplify.

OpenStudy (anonymous):

Or maybe you meant: \[\frac{d}{dx}(x^n)=nx^{n-1}\]. Whoops lol. If so, disregard above^^. (even though it works)

OpenStudy (anonymous):

d/dx (f(x))=f'(x)=lim h->0 f(x+h)=f(x) / h f(x) = x^3 + 4x - 2 f(x+h)+ (x+h)^3 + 4(x+h) - 2 x^3 +3x^2h+3xh^2+h^3+4x+4h-2 d/dx (x^3+4x-2) = lim h-> 0 (x^3+3x^2h+3xh^2+h^3+4x+4h-2)-(x^3+4x-2) / h = lim h->0 3x^2h + 3xh^2 + h^3 + 4h / h = lim h-> 0 3x^2 + 3xh + 4 = 3x^2+4